OpenStudy (anonymous):

Pls help solve 12y=9y^2+3

OpenStudy (anonymous):

rearrange first 12y=9y^2+3 0=9y^2-12y+3 0=3*(3y^2-4y+1) 0=3y^2-4y+1 0=(y-1)(3y-1) either y=1 or y=1/3